Circular No. 7435

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 V1 (SOHO) D. A. Biesecker, Emergent Information Technologies, Inc., and God-

dard Space Flight Center, reports SOHO observations of another Kreutz sungrazing comet discovered by D. Hammer and Biesecker using the automated search algorithm developed by Biesecker. The comet, which showed a tail, was seen in both the C2 and C3 telescopes. Astrometry (measured by Hammer and Biesecker, reduced by B. G. Marsden) and orbital elements appear on MPEC 2000-L07.

1998 UT  α_{2000}  δ_{2000}
Nov. 4.146  14^{h}23.^{m}5  −17^{o}27′

COMET C/2000 K6 (SOHO) Biesecker also reports observations of another Kreutz sungrazing comet

found via the SOHO website by M. Oates, X. Leprette, and M. Boschat. The comet, which showed no tail, was visible in the C2 telescope only. Astrometry (measured by Hammer and Biesecker, reduced by Marsden) and orbital elements appear on MPEC 2000-L08.

2000 UT  α_{2000}  δ_{2000}
May 27.088  4^{h}17.^{m}9  +19^{o}42′

SUPERNOVAE 1998A, 1998S, AND 1999gn S. D. Van Dyk and B. Nelson, Infrared Processing and Analysis Cen-

ter, California Institute of Technology; and the Two Micron All Sky Survey (2MASS) team write: “During routine 2MASS operations at Mt. Hopkins and Cerro Tololo, the type-II supernovae 1998A, 1998S, and 1999gn were observed. We obtained near-infrared photometry as part of pipeline processing. However, complicated backgrounds, such as host-galaxy spiral arms, limit the accuracy of the point-spread-function-fitting photometry procedure in the pipeline. Therefore, DAOPHOT (cf. Stetson 1987, PASP 99, 191) PSF-fitting photometry was employed to obtain the following magnitudes: SN 1998A in IC 2627, 1998 Apr. 27.13 UT, J (1.25 μm) = 15.40 ± 0.06, H (1.65 μm) = 14.94 ± 0.09, K_{s} (2.17 μm) = 14.94 ± 0.16. SN 1998S in NGC 3877, 1998 Dec. 24.48, J = 15.75 ± 0.10, H = 14.47 ± 0.06, K_{s} = 13.35 ± 0.04. SN 1999gn in NGC 4303 (M61), 2000 Feb. 25.27, J = 14.69 ± 0.04, H = 14.40 ± 0.06, K_{s} = 14.06 ± 0.07.”

2000 June 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green